Halloween Costumes Unlimited is considering a new 3-year store expansion project that requires an initial fixed asset investment of $3.3 million. The fixed asset falls into the 3-year MACRS class (MACRS Table) and will have a market value of $260,400 after 3 years. The project requires an initial investment in net working capital of $372,000. The project is estimated to generate $2,976,000 in annual sales, with costs of $1,190,400. The tax rate is 33 percent and the required return on the project is 9 percent. (Do not round your intermediate calculations.)

    

Required:
(a) What is the project's year 0 net cash flow?
  (Click to select)   -1,403,384   -3,488,400   -1,481,350   -3,672,000   -3,304,800

   

(b) What is the project's year 1 net cash flow?
  (Click to select)   1,637,281   1,403,384   1,559,316   1,481,350   1,715,247

  

(c) What is the project's year 2 net cash flow?
  (Click to select)   1,764,433   1,680,412   1,596,392   1,637,281   1,403,384

  

(d) What is the project's year 3 net cash flow?
  (Click to select)   1,637,281   1,984,796   1,786,316   1,885,556   2,084,036

  

(e) What is the NPV?
  (Click to select)   7,619,179   593,276   740,839   705,560.97   670,283
Answer #1

1.
=-3.3*10^6-372000=-3672000

2.
=(2976000-1190400-3.3*10^6*33.33%)*(1-33%)+3.3*10^6*33.33%=1559315.7

3.
=(2976000-1190400-3.3*10^6*44.45%)*(1-33%)+3.3*10^6*44.45%=1680412.5

4.
=(2976000-1190400-3.3*10^6*14.81%)*(1-33%)+3.3*10^6*14.81%+372000+260400-(260400-3.3*10^6*7.41%)*33%=1984795.8

5.
=-3672000+1559315.7/1.09+1680412.5/1.09^2+1984795.8/1.09^3
=705560.97